OpenStudy (anonymous):

simplify 10/8-5i i need someone to explain and help me work it out.i will reward a medal

OpenStudy (anonymous):

@Azureilai

OpenStudy (azureilai):

\[\frac{ 10 }{ 8-5i}\] You can seperate it into this \[\frac{ 10 }{ 8 }-\frac{ 10 }{ 5i }\] Then you simplify 10/8 And for 10/5i, it is the same as \[\frac{ 10 }{ \sqrt{-5} }\] You can multiply that by the irrational denominator so \[\frac{ 10 }{ \sqrt{-5} } \times \frac{ \sqrt{-5} }{ \sqrt{-5} }\] It will become \[\frac{ 10*\sqrt{-5} }{ -5 }\] I can only get to that far because I am not sure how to simplify the top. I think you can still simplify it, but I am not sure.

OpenStudy (anonymous):

ok thanks man i really apreciate it.and i got it man you multiply it by 10 and you get -50 so it would be 80-50i/39

OpenStudy (azureilai):

No problem, you might want to double check just in case

OpenStudy (azureilai):

http://www.wtamu.edu/academic/anns/mps/math/mathlab/int_algebra/int_alg_tut41_rationalize.htm This website might also help

OpenStudy (anonymous):

thanks man ur AWSOME!!!!!!!!